OpenStudy (anonymous):

How do you solve 3 over b^3 + 3 over b^2 equals 6 over b?

OpenStudy (anonymous):

please use the equation editor so that i may understand your question clearly. :)

OpenStudy (anonymous):

3/(b^3+3)/b^2=6/b? Then it becomes 3b^2/(b^3+3)=6/b ==> 3b^3/(b^3+3)=6 ==> 3b^3=6(b^3+3) ==> 3b^3=6b^3+18 ==> 6b^3-3b^3=-18 ==> 3b^3=-18 ==> :. b^3=-6 Hope this helps...

OpenStudy (anonymous):

3 3 6 --- + --- = -- b^3 b^2 b

OpenStudy (anonymous):

get the lcd, in this case is b^3 and multiple b^3 to the whole equation to get 3+3b=6b^2 6b^2-3b-3=0 2b^2-b-1 (2b+1)(b-1)=0 b=-1/2, b=1 hopefully i didn't have any mistakes there!

OpenStudy (anonymous):

How is it negative?

OpenStudy (anonymous):

(2b+1)(b-1)=0 equate each to 0 2b+1=0 2b=-1 b==1/2 b-1=0 b=1 :)

OpenStudy (anonymous):

i meant, b=-1/2 there. whoops!